Let X be a random variable with a normal distribution having a mean of 30 and a known standard deviation of 16. What is the probability that X is greater than 50?

A- 0.1056
B- 0.6057
C- 0.3944
D- 0.8944
Answer #1

Probability that X is greater than 50 = P(X>50) = 1- P(X<50) = 1-P(z<1.25) = 1- 0.8944 = 0.1056

where , Z score = ( X - mean)/std deviation = (50 - 30)/16

= 20/16 = 5/4 = 1.25

Correct option is A)